Now suppose the initial construction cost for a baseball park is $600 mln in year 2017, $300 mln in year 2018, and $300 mln in year 2019 . The useful economic life of the new ballpark is 20 years, and the appropriate discount rate is 10%. Assume that the expected operating profit is $150 mln a year.

Suppose the ballpark's construction starts at the end of 2017, lasts for three years (2017-2019), and it starts generating profit in 2020 (and works for 20 years starting in 2020).

What is the Net Present Value of the ballpark? (in mln)

Note 1: Discount the first year (2017) when calculating PV, since the construction starts at the END of that year.

Note 2:   Round your answer to the second decimal place: X.XX (mln)

Answer: -59.33

Need help on how they got this.

(B) Refer to the previous question. Is investing in the ballpark economically feasible?

True or False
Answer #1

a)

First we estimate the PV of Costs=PVC=-600*(P/F,10%,1)-300*(P/F,0.10,2)-300*(P/F,0.10,3)

(P/F,0.10,1)=1/(1+0.10)^1=0.909091

(P/F,0.10,2)=1/(1+0.10)^2=0.826446

(P/F,0.10,3)=1/(1+0.10)^3=0.751315

PV of Costs=PVC=-600*0.909091-300*0.826446-300*0.751315=-$1018.78 mln

Now estimate the PV of profits

PV of profits=PVR=150*(P/F,0.10,3)*(P/A,0.10,20)

PV of profits=PVR=150*0.751315*8.513564=$959.45 mln

NPV=PVC+PVF=-1018.78+959.45=-$59.33 mln

b)

NPV is negative. It means investing in ballpark is not economically feasible.